A spherical shaped vessel of 2 m diameter is 90 mm thick. Find the rate of heat leakage, if the temperature difference between the inner and outer surfaces is 220°C Thermal conductivity of the material of the sphere is 0.09 W/m°C.





1
Expert's answer
2022-01-10T09:09:43-0500

Solution.

"r_2=2m;"

"r_1=1.91m;"

"\Delta t=220^oC;"

"k=0.09W\/m^oC;"

"Q=\dfrac{\Delta t}{R};"

"R=\dfrac{r_2-r_1}{4\pi k r_1r_2};"

"R=\dfrac{2-1.91}{4\sdot3.14\sdot0.09\sdot2\sdot1.91}=\dfrac{0.09}{4.32}=0.021^oC/W;"

"Q=\dfrac{220}{0.021}=10476W=10.476kW;"

Answer: "Q=10.476kW."